Dina, Rami, Noura, and Jude were partners who shared profits and losses on a 4:2:2:2 basis, respectively. They were beginning to liquidate their business. At the start of the process, capital balances were as follows: Dina Capital $72,000 Rami capital 32,000 Noura capital 52,000 Jude capital 24,000 Who is the last to be paid if a pre-distribution plan is prepared?    A.   Dina    B.   Rami    C.   Jude    D.   Noura
Answer #1

Answer: option D; partner N

Step 1:

Sensitivity table

Partners

Capital / Ratio

Maximum loss absorption

D

72,000 / 0.4 =

180,000

R

32,000 / 0.2 =

160,000

N

52,000 / 0.2 =

260,000

J

24,000 / 0.2 =

120,000 Lowest

Therefore, if $120,000 is absorbed

D = 72,000 – (120,000 × 0.4) = 72,000 – 48,000 = 24,000

R = 32,000 – (120,000 × 0.2) = 32,000 – 24,000 = 8,000

N = 52,000 – (120,000 × 0.2) = 52,000 – 24,000 = 28,000

J = 24,000 – (120,000 × 0.2) = 24,000 – 24,000 = 0

Therefore, J would be paid first.

Step 1:

New ratio: D = 4/8 = 0.5, R = 2/8 = 0.25, N = 2/8 = 0.25

Sensitivity table

Partners

Capital / Ratio

Maximum loss absorption

D

24,000 / 0.5 =

48,000

R

8,000 / 0.25 =

32,000 Lowest

N

28,000 / 0.25 =

112,000

Therefore, if $32,000 is absorbed

D = 24,000 – (32,000 × 0.5) = 24,000 – 16,000 = 8,000

R = 8,000 – (32,000 × 0.25) = 8,000 – 8,000 = 0

N = 28,000 – (32,000 × 0.25) = 28,000 – 8,000 = 20,000

Therefore, R would be paid second.

Step 3:

New ratio: D = 4/6 = 0.67, N = 2/6 = 0.33

Sensitivity table

Partners

Capital / Ratio

Maximum loss absorption

D

8,000 / 0.67 =

11,940 Lowest

N

20,000 / 0.33 =

60,606

Therefore, if $11,940 is absorbed

D would be paid third and N would be paid last.